An open box is to be made from a rectangular piece of tin by cutting two inch squares out of the corners and folding up the sides. The volume of the box will be 100 cubic inches. Find the dimensions of the rectangular piece of tin.
Answer by checkley77(12844)

Let L= the length or the base
Let w= the width of the base.
AREA=H*LW
100=2*LW
LW=100/2
LW=50
ANY SET OF L TIMES W VALUES THAT = 50 ARE ANSWERS FOR THE BASE OF THE BOX.
5*10=50
PROOF:
100=2*5*10
100=100
NOW ADD 4 INCHES TO EACH OF THE BASE MEASUREMENTS FOR THE CORNER CUTS
TO FIND THE ORIGINAL SIZE OF THE PIECE OF TIN.
(5+4)=9 INCHES
(10+4)=14 INCHES
THESE COULD BE THE ORIGINAL MEASUREMENTS OF THE TIN.